# Frequency Masking for Universal Deepfake Detection

### 12 Jan 2024 v1

@@ Abstract
We study universal deepfake detection. Our goal is to detect synthetic images from a range of generative AI approaches, particularly from emerging ones which are unseen during training of the deepfake detector. Universal deepfake detection requires outstanding generalization capability. Motivated by recently proposed masked image modeling which has demonstrated excellent generalization in self-supervised pre-training, we make the first attempt to explore masked image modeling for universal deepfake detection. We study spatial and frequency domain masking in training deepfake detectors. Based on empirical analysis, we propose a novel deepfake detector via frequency masking. Our focus on frequency domain is different from the majority, which primarily target spatial domain detection. Our comparative analyses reveal substantial performance gains over existing methods. Code and models are publicly available [^1] https://github.com/chandlerbing65nm/FakeDetection.git.
@@ 1 Introduction
In our study, we analyze both spatial and frequency masking for universal deepfake detection. Our results suggest that frequency masking is more effective than spatial masking in generalizing deepfake detection for different generative AI approaches. Our finding is consistent with a recent study by Corvi et al. [14], which discovers frequency artifacts in GAN and diffusion model-based synthetic images. Different from [14], our main contribution is a new training method to improve detection accuracy via frequency masking. We remark that most existing detectors focus on spatial domain [2, 7]. Our contributions are summarized as follows:
@@ 2.2 Frequency Domain Masking
Given the selected frequency band $[u_{\text{start}},u_{\text{end}}]\times[v_{\text{start}},v_{\text{end}}]$, the number of frequencies to be nullified $N$ is calculated: $N=\lceil r\times(u_{\text{end}}-u_{\text{start}})\times(v_{\text{end}}-v_{\text{start}})\rceil$. Frequencies within this region are set to zero, resulting in a masked frequency representation $M(u,v)$:
@@
Table 3: Generalization results showcasing the effectiveness of our frequency-based masking technique. We compare the Average Precision (AP) with Wang et al. [7] and Gragnaniello et al. [2], which are SOTA according to a recent study [15]. When integrated with these SOTA methods, our approach consistently improves mAP by significant amounts, highlighting the utility of frequency masking in enhancing universal deepfake detection.
@@ 3.1 Comparison of Mask Types
For masking types, as depicted in Figure 2, we observe a distinct hierarchy in terms of mean average precision (mAP). Pixel masking exhibits the lowest performance with an mAP of 75.12%. In contrast, patch masking with random 8x8 blocks of pixels as patches shows a notable improvement with an mAP of 86.09%. However, the most interesting result is the superior performance of frequency-based masking, which reaches an mAP of 88.22%. This clearly indicates that frequency-based masking captures more generalizable features compared to pixel and patch masking, thereby substantiating its efficacy in universal deepfake detection task.
@@ 3.4 Comparison with State-Of-The-Art
As shown in Table 3, ‘+Ours’ approach, which incorporates $15\%$ random masking in ’all’ frequency band, demonstrates a consistent improvement in performance when integrated with existing SOTA methods. For instance, when combined with Wang et al.’s [7] method, our frequency-based masking technique yields an improvement of $3.37\%$ and $2.36\%$ in mean average precision (mAP) for the Blur+JPEG (0.5) and Blur+JPEG (0.1) variants, respectively. Notably, the integration of our method with Gragnaniello et al.’s approach [2] led to a significant improvement of $3.01\%$. This is a compelling testament to the efficacy of frequency masking, as it enables the model to learn more generalizable features for universal deepfake detection that are potentially overlooked by Wang et al.’s [7] approach.
Moreover, our approach also shines when coupled with Ojha et al.’s [1] method which is based on visual encoder of CLIP [16], manifesting in a significant mAP boost of $0.68\%$. This shows the robustness of our frequency masking strategy across different SOTA methods. By strategically masking random frequencies, our model learns more generalizable features. Thus, these results shows our approach as valuable asset in the pursuit of more precise universal deepfake detection.
@@ 4 Conclusion
Motivated by recent promising results of masked image modeling, we proposed a frequency-based masking strategy tailored for universal detection of deepfakes. Our method enables deepfake detector to learn generalizable features in the frequency domain. Our experiments, conducted using samples with a range of generative models, empirically demonstrate the efficiency of our proposed approach. The superior performance of our method highlights the potential of masked image modeling and frequency domain approach for the challenging problem of universal deepfake detection.

### 17 Jan 2024 v3

@@ Abstract
We study universal deepfake detection. Our goal is to detect synthetic images from a range of generative AI approaches, particularly from emerging ones which are unseen during training of the deepfake detector. Universal deepfake detection requires outstanding generalization capability. Motivated by recently proposed masked image modeling which has demonstrated excellent generalization in self-supervised pre-training, we make the first attempt to explore masked image modeling for universal deepfake detection. We study spatial and frequency domain masking in training deepfake detectors. Based on empirical analysis, we propose a novel deepfake detector via frequency masking. Our focus on frequency domain is different from the majority, which primarily target spatial domain detection. Our comparative analyses reveal substantial performance gains over existing methods. Code and models are publicly available [^1] https://github.com/chandlerbing65nm/FakeImageDetection.git.
@@ 1 Introduction
In our study, we analyze both spatial and frequency masking for universal deepfake detection. Our results suggest that frequency masking is more effective than spatial masking in generalizing deepfake detection for different generative AI approaches. Our finding is consistent with a recent study by Corvi et al. [14], which discovers frequency artifacts in GAN and diffusion model-based synthetic images. Different from [14], our main contribution is a new training method to improve detection accuracy via frequency masking. We remark that most existing detectors focus on spatial domain [1, 2, 7]. Our contributions are summarized as follows:
@@ 2.2 Frequency Domain Masking
Given the selected frequency band $[u_{\text{start}},u_{\text{end}}]\times[v_{\text{start}},v_{\text{end}}]$, the frequencies to be nullified $N$ is calculated: $N=\lceil r\times(u_{\text{end}}-u_{\text{start}})\times(v_{\text{end}}-v_{\text{start}})\rceil$. Frequencies within this region are set to zero, resulting in a masked frequency representation $M(u,v)$:
@@
Table 3: Generalization results showcasing the effectiveness of our frequency-based masking technique. We compare the Average Precision (AP) with Wang et al. [7], Gragnaniello et al. [2], and Ojha et al. [1], which are SOTA according to a recent study [15]. When integrated with these SOTA methods, our approach consistently improves mAP by significant amounts, highlighting the utility of frequency masking in enhancing universal deepfake detection.
@@ 3.1 Comparison of Mask Types
For masking types, as depicted in Figure 2, we observe a distinct hierarchy in terms of mean average precision (mAP). Pixel masking exhibits the lowest performance with an mAP of 75.12%. In contrast, patch masking with random 8x8 blocks of pixels as patches shows a notable improvement with an mAP of 86.09%. However, the most interesting result is the superior performance of frequency-based masking, which reaches an mAP of 88.22%. This clearly indicates that frequency-based masking captures more generalizable features compared to pixel and patch masking, thereby substantiating its effectiveness in universal deepfake detection task.
@@ 3.4 Comparison with State-Of-The-Art
As shown in Table 3, ‘+ Ours’ approach, incorporating $15\%$ random masking across ‘all’ frequency bands, consistently enhances performance when integrated with existing state-of-the-art (SOTA) methods. Specifically, the combination of our frequency-based masking technique with Wang et al.’s method [7] results in an increase of $3.37\%$ and $2.36\%$ in mean average precision (mAP) for the Blur+JPEG (0.5) and Blur+JPEG (0.1) variants, respectively. Furthermore, integrating our method with that of Gragnaniello et al. [2] yields a notable improvement of $3.01\%$. These are testaments to the effectiveness of frequency masking, as it enables the model to learn more generalizable features for universal deepfake detection that are potentially overlooked by Wang et al.’s [7] and Gragnaniello et al.’s [2] approaches.
Moreover, our approach also shines when coupled with Ojha et al.’s [1] method which is based on visual encoder of CLIP [16], manifesting in a significant mAP boost of $0.68\%$. This further affirms the adaptability of our approach across different backbones for universal deepfake detection. By strategically masking random frequencies, our model learns more generalizable features. Thus, these results shows our approach as valuable asset in the pursuit of more precise universal deepfake detection.
@@ 4 Conclusion
Motivated by recent promising results of masked image modeling, we proposed a frequency-based masking strategy tailored for universal detection of deepfakes. Our method enables deepfake detector to learn generalizable features in the frequency domain. Our experiments, conducted using samples with a range of generative models, empirically demonstrate the effectiveness of our proposed approach. The superior performance of our method highlights the potential of masked image modeling and frequency domain approach for the challenging problem of universal deepfake detection.
5 Acknowledgement
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programmes (AISG Award No.: AISG2-TC-2022-007) and SUTD project PIE-SGP-AI-2018-01. This research work is also supported by the Agency for Science, Technology and Research (A*STAR) under its MTC Programmatic Funds (Grant No. M23L7b0021). This material is based on the research/work support in part by the Changi General Hospital and Singapore University of Technology and Design, under the HealthTech Innovation Fund (HTIF Award No. CGH-SUTD-2021-004). We thank anonymous reviewers for their insightful feedback.